OpenStudy (anonymous):

Simplify 10P5 Simplify 10C3

OpenStudy (yanasidlinskiy):

Do you have a TI-84 calculator?

OpenStudy (anonymous):

No, only a simple one.

OpenStudy (yanasidlinskiy):

Oh, okay. Well there's a website calculator. Would you like to know? Or do you want to know how?

OpenStudy (anonymous):

Yes, I would like to know!

OpenStudy (yanasidlinskiy):

https://www.mathway.com/ You can either use that website or....

OpenStudy (anonymous):

Thank you very much!

OpenStudy (yanasidlinskiy):

Would you like to know how to do it by hand?

OpenStudy (yanasidlinskiy):

Permutation Formula: nPk=n!/(n-k)!

OpenStudy (yanasidlinskiy):

10!/(10-5)! Calculate the numerator: 10! 10!= 10*9*8*7*6*5*4*3*2*1 10!= 3,68,800 Calculate the denominator: (n-r)! = (10-5)! (10-5)! = 5! 5!= 5*4*3*2*1 5!= 120 10P5= 3,628,800/120 10P5= 30,240

OpenStudy (anonymous):

Can I ask where you got (n-r)! from?

OpenStudy (yanasidlinskiy):

Sorry *k It's the formula.

OpenStudy (yanasidlinskiy):

Permutation = nPr = n! / ( n - r )! Combination = nCr = nPr / r! where, n, r are non negative integers and r <= n. r is the size of each permutation n is the size of the set from which elements are permuted. !is the factorial operator.

OpenStudy (yanasidlinskiy):

Does that make sense?

OpenStudy (anonymous):

Yes! Thanks!

OpenStudy (yanasidlinskiy):

You're welcome:) Glad I could help:)